Compute RNOA with Disaggregation

$ millions May. 10, 2020 Feb. 20, 2019
Operating assets $42,225 $40,934
Nonoperating assets 1,778 3,595
Total assets $44,003 $44,529
Operating liabilities $16,679 $16,047
Nonoperating liabilities 29,202 27,028
Total liabilities $45,881 $43,075
Net sales $108,203
Operating expense before tax 92,673
Net operating profit before tax (NOPBT) 15,530
Other expense 974
Income before tax 14,556
Tax expense 3,435
Net income $11,121


Round all answers to two decimal places (ex: 0.12345 = 12.35%)

a. Compute return on net operating assets (RNOA).

RNOA = 48.04 (incorrect)

b. Disaggregate RNOA into components of profitability and productivity and show that the product of the two components equals RNOA. Assume a statutory tax rate of 22%.

NOPM = 11.19% (incorrect)

NOAT = 4.29 (correct)
Answer #1
Part A Solution
Return on net operating asset (RNOA) = Net operating profit after tax/Net operating asset
Current tax rate: Tax expense/ Income before tax
.=3435/14556 = 23.6%
Net operating profit before tax (NOPBT)= 15,530
Net operating profit after tax (NOPAT)= 15,530 * (1-.236) = 11,864.92
Avg. operating asset = (42,225 + 40,934)/2 = 41,579.5
Avg. operating liability = (16,679 + 16,047)/2 = 16,363
Net operating asset = 41,579.5 - 16,363 = 25,216.5
RNOA = 11,864.92/25,216.5 = 47.05%
Part B Solution
RNOA= NOPAT/Avg. NOA = (NOPAT/Sales) * (Sales/Avg. NOA)
Net operating profit before tax (NOPBT)= 15,530
New tax rate = 22%
Net operating profit after tax (NOPAT)= 15,530 * (1-.22) = 12,113.4
Sales = 108,203
Net operating asset = 41,579.5 - 16,363 = 25,216.5
Net operating profit margin (NOPM): Net operating profit after tax/ Sales
=12,113.4/108,203 = 11.19%
Net operating asset turnover (NOAT): Sales/ net operating asset
=108,203/25,216.5 = 4.29